OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1

OpenStudy (anonymous):

{2*x + y = 1, 4*x + 2*y = -1} Divide each side of the second equation by 2. {2*x + y = 1, 2*x + y = -1/2} 1 does not equal -1/2 There is no solution.

hero (hero):

2x + y = 1 2x + y = -1/2 False